Circular No. 8920

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2007 X5–X15 (SOHO) Further to IAUC 8919, additional near-sun comets have been found on

SOHO website images; all were Kreutz sungrazers except for Meyer-group objects C/2007 X7 and C/2007 X14, which were both very small and stellar in appearance (mag ∼ 7). C/2007 X5 was stellar in appearance and small (mag ∼ 6.5) in C3 images, and slightly diffuse with no tail in C2 images. K. Battams describes C/2007 X6 as extremely faint (mag ∼ 8.5), diffuse, and small. C/2007 X8, C/2007 X9 (which was also found by T. Hoffman), and C/2007 X12 were tiny and stellar in appearance in C3 images (reaching mag ∼ 7, except C/2007 X12, which reached mag ∼ 7.5); in C2 images, they appeared diffuse, with C/2007 X8 appearing a little elongated and the other two objects showing no tail. C/2007 X10 was teardrop-shaped and very bright (mag ∼ 3) in C3 images, and showed a broad head and a thin tail in C2 images. C/2007 X11, which was also found by H. Su, was very diffuse and faint (mag ∼ 8). C/2007 X13 was very diffuse, elongated, and reached mag ∼ 7.5. C/2007 X15 was small and diffuse (mag ∼ 7.5). New finder codes: MK = M. Kusiak; PC = P. Collison.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 X5 Dec. 7.679 16^{h}53.^{m}4 −26^{o}46′ C3/2 HS 2008-B36
C/2007 X6  10.314 17 09.1 −24 44 C2  MK 2008-B45
C/2007 X7  10.381 17 06.0 −21 50 C2  HS 2008-B45
C/2007 X8  10.638 17 10.0 −25 31 C3/2 TH 2008-B45
C/2007 X9  12.179 17 17.2 −25 26 C3/2 RM 2008-B45
C/2007 X10  12.488 17 13.5 −29 42 C3/2 TH 2008-B46
C/2007 X11 12.564 17 19.5 −24 55 C2 RK 2008-B46
C/2007 X12  13.738 17 24.5 −25 26 C3/2 RM 2008-B46
C/2007 X13  14.088 17 26.2 −25 04 C2  PC 2008-B46
C/2007 X14  14.163 17 22.9 −22 10 C2  TH 2008-B48
C/2007 X15  15.479 17 34.8 −25 02 C2  HS 2008-B48

COMET 17P/HOLMES Visual total-magnitude and coma-diameter estimates by A. Pereira,

Cabo da Roca, Portugal (5×22 monocular): 2007 Dec. 26.81 UT, 3.5, ∼ 68′; 29.88, 3.6, > 65′; 2008 Jan. 1.08, 3.6, ∼ 60′; 4.03, 3.5, ∼ 78′; 12.03, 3.8, > 65′; 26.81, 3.8, ∼ 80′; 28.92, 4.0, ∼ 84′; 31.95, 4.1, > 70′; Feb. 2.84, 4.3, ∼ 70′; 4.87, 4.5, ∼ 80′; 7.97, 4.6, ∼ 85′; 10.02, 4.6, ∼ 90′.

2008 February 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green